OpenStudy (anonymous):

Find the sum of the geometric sequence. 3/2, 3/8, 3/32, 3/128, 3/512

OpenStudy (anonymous):

i thin it is 341

OpenStudy (anonymous):

please help anybody

OpenStudy (anonymous):

@stamp

OpenStudy (stamp):

sum of a geometric series: http://www.mathsisfun.com/algebra/sequences-sums-geometric.html Sn = a(1 - r^n)/(1 - r) r = 1/4, a = 3/2, n = 5 \[S_5=(3/2)(1-(1/4)^5)/(1-(1/4))\] Sum =1023/512 = http://www.wolframalpha.com/input/?i=%283%2F2%29%281-%281%2F4%29%5E5%29%2F%281-%281%2F4%29%29

OpenStudy (anonymous):

thankyou